If five times David's age plus 6 years is the same as the square of his age, how old is david?
Answer by stanbon(75887)

If five times David's age plus 6 years is the same as the square of his age, how old is david?
----------
Equation::
5d + 6 = d^2
--------
d^2 - 5d - 6 = 0
--------
(d-6)(x+1) = 0
----
d = 6 yrs (David's age now)
-----------
Cheers,
Stan H.
-----------